You invested $6000 at 4.5% and 5.5% simple interest. During the first year, the two accounts earned $305. How much did you invest in each fund? (note: the 5.5% account is more risky.)
Word problems aren't usually that difficult, but for whatever reason, this one has confused me. I don't know why...I think it's hard for me to keep track of as it progresses. Could you please help me? Thanks so much!

Answer by checkley77(12844)

.055x+.045(6,000-x)=305
.055x+270-.045x=305
.-1x=305-270
.01x=35
x=35/.01
x=$3,500 is the amount invested @ 5.5%
6,000-3,500=$2,500 is the amount invested @ 4.5%
Proof:
.055*3,500+.045*2,500=305
192.50+112.50=305
305=306